How many particles of H2SO4  are needed to react completely with 1.89 moles of H2O ?


1
Expert's answer
2021-04-14T06:13:27-0400

H20+SO3 -> H2SO4

Mole ratio = 1:1

Therefore 1.89 moles of H2O = 1.89 moles of H2SO4

1 mole = 6×1023 particles

= 1.89 × 6× 1023

= 1.134 × 1024 particles